# On some aspects of local thermodynamical formalism

### v2

@@ 2.5 Translocal entropy for higher-dimensional maps
$h_{\text{\tiny loc}}(f)=\lim_{\varepsilon\to 0}h_{\text{\tiny loc}}(f,\varepsilon)\quad\text{ for }\quad h_{\text{\tiny loc}}(f,\varepsilon)=\limsup_{n\to\infty}\frac{1}{n}\sup_{x\in X}\log S(n,\varepsilon,B_{n}(z,\varepsilon)).$ (5)
In fact, Buzzi uses $(n,\varepsilon)$-spanning sets rather than $(n,\varepsilon)$-separating sets, but that is of no consequence. Also, by taking the supremum over $x\in X$, $h_{\text{\tiny loc}}(f)$ becomes a quantity independent of the point in space. Buzzi shows that $h_{\text{\tiny loc}}(f)\leq\frac{\dim M}{r}\log\inf_{n}\root n \of{\|Df^{n}\|_{\infty}}$ if $f:M\to M$ is a $C^{r}$-smooth map on a manifold $M$, and hence $h_{\text{\tiny loc}}(f)=0$ for $C^{\infty}$ maps. But if we adjust Example 2.5 to $(r,\phi)=(r,3\phi\pmod{2\pi})$, then the point 00 gives local entropy $\log 3$.
@@ 3.1 Brin-Katok local measure entropy
For a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$ and $f$-invariant Borel probability measure $\mu$, Brin & Katok [11] introduced notions of local lower (resp. local upper) measure-theoretic entropy w.r.t. $\mu$ as follows
@@
Theorem 3.1 ([11])
For a continuous map $F:X\to X$ on a compact metric space $(X,d)$ and Borel $F$-invariant probability measure $\mu$, the equality ${\underline{h}}_{\mu}(f,x)={\overline{h}}_{\mu}(f,x):=h_{\mu}(f,x)$ holds for $\mu$-a.e. $x\in X$ Moreover, $h_{\mu}(f,x)$ is F-invariant and
[…]
In addition, for an ergodic measure $\mu$ the equality $\overline{h}_{\mu}(f,x)=h_{\mu}(f)$ holds for $\mu$-a.e. $x\in X$. A variational principle for the Brin-Katok entropy was proved by Feng & Huang [18]:
$h_{\text{\tiny top}}(f,K)=\sup\{\underline{h}_{\mu}(f):\mu\text{ is ergodic and $f$-invariant and }\mu(K)=1\},$ (7)
where $h_{\text{\tiny top}}(f,K)$ is as in (1) for a compact set $K$.
Ma & Wen [27] noticed relationship between topological entropy restricted to a set $Z\subset X$ and local measure entropies on $Z$.
Theorem 3.2 ([27])
@@
    If ${\underline{h}}_{\mu}(f,x)\leq s$ for all $x\in X$, then $h_{\text{\tiny top}}(f,Z)\leq s$,
@@
    ${\overline{h}}_{\mu}(f,x)\geq s$ for all $x\in X$ and $\mu(Z)>0$, then $h_{\text{\tiny top}}(f,Z)\geq s$.
In the next section we provide a generalization of Theorem 3.2.
3.2 Topological pressure and Carathéodory structures
We recall the notion of topological pressure, in the spirit of Carathéodory structures elaborated by Pesin [32], for a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$, and restricted to a subset $Z\subset X$. Fix a continuous map (called potential) $\phi:X\to{\mathbb{R}}$, $N\in{\mathbb{N}}$, $r>0$ and $s\geq 0$, define
@@ 3.2 Topological pressure and Carathéodory structures
where the infimum is taken over all finite or countable coverings of the $Z$ by Bowen balls $\{B_{n_{j}}(x_{j},r)\}_{j\in J}$ with $n_{j}\geq N.$ Here, the Bowen balls are indexed by the elements of a finite or countable set $J$ Using standard arguments (see [32]) one can easily show that for any $N\in{\mathbb{N}}$, the inequality $M_{Z}(s,r,\phi,N+1)\geq M_{Z}(s,r,\phi,N)$ holds. Therefore, there exists a limit
@@
The graph of the function $s\mapsto M_{Z}(s,r,\phi)$ is very similar to the graph of $s$-Hausdorff measure function, i.e., there exists a unique critical parameter $s_{0}$, where the function $s\mapsto M_{Z}(s,r,\phi)$ drops from $\infty$ to 00. Thus, we can define
@@
Fix $r_{1}\leq r_{2}$ and consider a cover $\{B_{n_{j}}(x_{j},r_{1})\}_{j\in J}$ of $Z$, with $n_{j}\geq N$. Then $\{B_{n_{j}}(x_{j},r_{2})\}_{j\in J}$ is a covering of $Z$, so $M_{Z}(s,r_{1},\phi)\geq M_{Z}(s,r_{2},\phi)$. Therefore
@@ 3.3 Local measure-theoretic pressures
Notice that local measure-theoretic pressures, calculated with respect to the potential $\phi\equiv 0$, coincide with local measure-theoretic Brin-Katok entropies. Moreover, local measure-theoretic pressures and topological pressure are related. We have the following partial variational principle, extending Ma & Wen’s result Theorem 3.2.
@@
Given a continuous map $f:X\to X$ on a compact metric space, a Borel subset $Z\subset X$ and a Borel probability measure $\mu$ on $X$, we have for any $s\geq 0$ and continuous $\phi:X\to{\mathbb{R}}$:
@@
$Z_{k}:=\left\{x\in Z:\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\leq s+\varepsilon\text{ for all }r\in(0,\frac{1}{3k})\right\}.$
@@
$\mu(B_{n_{i}(x_{i})}(x_{i},r)))\geq\exp\left(-(s+\varepsilon)n_{i}(x_{i})+\sum_{m=0}^{{n_{i}(x_{i})}-1}\phi(f^{m}(x_{i}))\right).$
@@
$V_{k}:=\left\{x\in Z:\liminf_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\geq s-\varepsilon,\text{ for every }r\in(0,\frac{1}{k})\right\}.$
@@
$V_{k_{0},N}:=\left\{x\in Z:\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\geq s-\varepsilon,\text{ for all }n\geq N,r\in\left(0,\frac{1}{k_{0}}\right)\right\}.$
@@
Passing to the limit $N\to\infty$, we get $M_{Z}(s-\varepsilon,r,\phi)>0$. This means that $M_{Z}(r,\phi)>s-\varepsilon.$ Finally passing to the limit $r\to 0$ and taking into account that $\varepsilon$ is arbitrary small, we obtain $P_{Z}(f,\phi)\geq s$. The proof is complete. $\Box$
@@ 3.4 The translocal version of measure-theoretic local pressure
We suggest for translocal version for the upper measure-theoretic local pressure ${\overline{P}}_{\mu}(f,\phi,z)$ from (8) is the following:
$\overline{P}_{\mu,\omega}(f,\phi,z):=\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(z))-\log\mu(B(z,e^{-\omega n}))\right),$ (9)
@@
$\underline{P}_{\mu,\omega}(f,\phi,z):=\liminf_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(z))-\log\mu(B(z,e^{-\omega n}))\right),$ (10)
@@
where the infimum is taken over all finite or countable coverings of the $Z$ by metric balls $\{B(x_{j},e^{-\omega n_{j}})\}_{j\in J}$ with $n_{j}\geq N.$ Let
@@
The graph of the function $s\mapsto\overline{M}_{f,Z,\omega}(s,r,\phi)$ has a unique critical parameter $s_{0}$, where the function $s\mapsto\overline{M}_{f,Z,\omega}(s,r,\phi)$ drops from $\infty$ to 00. The graph of the function $s\mapsto\underline{M}_{f,Z,\omega}(s,r,\phi)$ has a unique critical parameter $s_{1}$, where the function $s\mapsto\underline{M}_{f,Z,\omega}(s,r,\phi)$ drops from $\infty$ to 00.
@@
A metric space X is called boundedly compact if all bounded closed subsets of X are compact. In particular ${\mathbb{R}}^{n}$ and Riemannian manifolds (see [22, p.9]) are boundedly compact.
@@
(Vitali Covering Lemma, Theorem 2.1 in [28]). Let X be a boundedly compact metric space and a family of closed balls =${\mathcal{B}}=\{\overline{B(x,r)}:x\in X,r>0\}$ in X such that
@@
$\bigcup_{B(x,r)\in{\mathcal{B}}}\subset\bigcup_{i\in I}\overline{B(x_{i},5\cdot r_{i})}.$
@@
Given a continuous map $f:X\to X$ on a compact metric space, a Borel subset $Z\subset X$ and a Borel probability measure $\mu$ on $X$, we have for any $s\geq 0$ and continuous $\phi:X\to{\mathbb{R}}$:
@@
2) If $\mu(Z)>0$ and ${\underline{P}}_{\mu}(f,\phi,x)\geq s$ or all $x\in Z$, then $\underline{P}_{Z,\omega}(f,\phi)\geq s$.
@@
[…]
@@
${\mathcal{C}}:=\left\{B_{n_{j}}(y_{j},e^{-\omega n_{j}}):y_{j}\in V_{N_{0}},~{}n_{j}\geq N_{0},~{}B(y_{j},e^{-\omega n_{j}})\cap V_{N_{0}}\neq\emptyset,j\in J\right\}.$
@@
[…]
@@ 4 Discussion on local entropy
In the literature on dynamical systems there are several attempts to define local entropy. Bowen [9] introduced the notion of $\varepsilon$-entropy to get an upper estimate of the difference between $h_{\mu}(f)$ and the entropy of a partition with diameter less than $\varepsilon$. Another notion of local entropy (called conditional topological entropy) was given by Misiurewicz [29], who proved that vanishing of his local entropy implies existence of maximal measure. Finding (uniqueness of) measures of maximal entropy was also the reason for Newhouse [30] and Buzzi & Ruette [13, 15] to define a version the local entropy $h_{\text{\tiny loc}}(f)$, see (5). Buzzi & Ruette [15] proved that for $C^{1}$ maps $f:[0,1]\to[0,1]$ with critical set $Crit(g)=\{x\in[0,1]:f|_{U}\text{ is not monotone on each neighborhood }U\owns x\}$, if $h_{\text{\tiny top}}(g)>h_{\text{\tiny top}}(g,Crit(g))+h_{\text{\tiny loc}}(f)$, then $f$ admits a measure of maximal entropy.
In 1993, Blanchard [6] introduced a notion of entropy pairs, in order to localize “where” in the system entropy is generated. This was generalized to entropy $n$-tuples and entropy sets by Huang & Ye [24] and Blanchard & Huang [7].

### v3

@@ 2.5 Translocal entropy for higher-dimensional maps
$h_{\text{\tiny loc}}(f)=\lim_{\varepsilon\to 0}h_{\text{\tiny loc}}(f,\varepsilon)\quad\text{ for }\quad h_{\text{\tiny loc}}(f,\varepsilon)=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\sup_{x\in X}\log r(n,\delta,B_{n}(x,\varepsilon)).$ (6)
In fact, Buzzi uses $(n,\delta)$-spanning sets (the minimal cardinality of which is denoted by $r$ in (6)) instead of $(n,\delta)$-separating sets, but that is of no consequence. Also, by taking the supremum over $x\in X$, $h_{\text{\tiny loc}}(f)$ becomes a quantity independent of the point in space. Buzzi shows that $h_{\text{\tiny loc}}(f)\leq\frac{\dim M}{r}\log\inf_{n}\root n \of{\|Df^{n}\|_{\infty}}$ if $f:M\to M$ is a $C^{r}$-smooth map on a manifold $M$, and hence $h_{\text{\tiny loc}}(f)=0$ for $C^{\infty}$ maps. But if we adjust Example 2.5 to $f(r,\phi)=(r,3\phi\pmod{2\pi})$, then the point $(0,0)$ gives local entropy $\log 3$.
@@ 3.1 Brin-Katok local measure entropy
For a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$ and $f$-invariant Borel probability measure $\mu$, Brin & Katok [12] introduced notions of local lower and local upper measure-theoretic entropy w.r.t. $\mu$ as follows
@@
Theorem 3.1 (Main Theorem and Corollary of [12])
For a continuous map $f:X\to X$ on a compact metric space $(X,d)$ and Borel $f$-invariant probability measure $\mu$, the equality ${\underline{h}}_{\mu}(f,x)={\overline{h}}_{\mu}(f,x):=h_{\mu}(f,x)$ holds for $\mu$-a.e. $x\in X$. Moreover, $h_{\mu}(f,x)$ is $f$-invariant and
$\int_{X}h_{\mu}(f,x)\,d\mu(x)=h_{\mu}(f).$
If $\mu$ is ergodic, then additionally $h_{\mu}(f,x)=h_{\mu}(f)$ for $\mu$-a.e. $x$.
Feng & Huang [21, Theorem 1.2(i)] showed in 2012 the following variational principle for the Brin-Katok local entropy. For a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$ and a compact non-empty subset $K\subset X$ they obtain
$h_{\text{\tiny top}}^{FH}(f,K)=\sup\{\underline{h}_{\mu}(f):\mu\text{ \color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}is a Borel probability measure and }\mu(K)=1\},$ (8)
where $\underline{h}_{\mu}(f)$ local lower entropy. However, here $h_{\text{\tiny top}}^{FH}(f,K)$ is an entropy-like quantity constructed in a way similar to the construction of $s$-Hausdorff measure and Hausdorff dimension, called Bowen topological entropy, see [21, Section 2.2]. This is based on a Carathéodory-like construction, but a bit more involved than the Bowen topological entropy used in Theorem 3.2 below, and which we discuss in the next section, generalized to pressure.
3.2 Topological pressure and Carathéodory structures
In 2008, Ma & Wen [32] noticed relationship between topological entropy restricted to a set $Z\subset X$ and local measure entropies on $Z$.
Theorem 3.2 (Theorem 1 in [32])
@@ 3.2 Topological pressure and Carathéodory structures
    If ${\underline{h}}_{\mu}(f,x)\leq s$ for all $x\in X$, then $h_{\text{\tiny top}}^{B}(f,Z)\leq s$.
@@
    ${\overline{h}}_{\mu}(f,x)\geq s$ for all $x\in X$ and $\mu(Z)>0$, then $h_{\text{\tiny top}}^{B}(f,Z)\geq s$.
The Bowen entropy $h_{\text{\tiny top}}^{B}$ used in this result is based on Carathéodory construction which we will discuss in this subsection at length, but for topological pressure, so with the inclusion of a potential function $\phi:X\to{\mathbb{R}}$. By setting $\phi\equiv 0$ in the topological pressure defined below, we get $h_{\text{\tiny top}}^{B}(f)$. Wang & Zhang [46] obtain a similar result for upper local entropies $\overline{h}_{\mu}(f,x)$ and packing topological entropies.
The aim of this subsection is to provide a generalization of Theorem 3.2 to topological pressure. Topological pressure, in the spirit of Carathéodory structures elaborated by Pesin [37], for a continuous map $f:X\to X$ defined on a compact metric space $(X,d)$, and restricted to a subset $Z\subset X$. Fix a continuous map (called potential) $\phi:X\to{\mathbb{R}}$, $N\in{\mathbb{N}}$, $r>0$ and $s\geq 0$, define
@@
where the infimum is over all finite or countable covers of $Z$ by Bowen balls $\{B_{n_{j}}(x_{j},r)\}_{j\in J}$ with $n_{j}\geq N$. Here, the Bowen balls are indexed by the elements of a finite or countable set $J$. Using standard arguments (see [37]) one can easily show that for any $N\in{\mathbb{N}}$, the inequality $M_{Z}(s,r,\phi,N+1)\geq M_{Z}(s,r,\phi,N)$ holds. Therefore, there exists a limit
@@
The graph of the function $s\mapsto M_{Z}(s,r,\phi)$ is very similar to the graph of $s$-Hausdorff measure function, i.e., there exists a unique critical parameter $s_{0}$, where the function $s\mapsto M_{Z}(s,r,\phi)$ drops from $\infty$ to 0. Thus, we can define
@@
Fix $r_{1}\leq r_{2}$ and consider a cover $\{B_{n_{j}}(x_{j},r_{1})\}_{j\in J}$ of $Z$, with $n_{j}\geq N$. Then $\{B_{n_{j}}(x_{j},r_{2})\}_{j\in J}$ is a cover of $Z$, so $M_{Z}(s,r_{1},\phi)\geq M_{Z}(s,r_{2},\phi)$. Therefore
@@ 3.3 Local measure-theoretic pressures
Notice that local measure-theoretic pressures, calculated with respect to the potential $\phi\equiv 0$, coincide with local measure-theoretic Brin-Katok entropies. Carvalho & Pérez [17] showed that if $\mu$ is ergodic, then ${\overline{P}}_{\mu}(f,\phi,x)={\underline{P}}_{\mu}(f,\phi,x)$ for $\mu$-a.e. $x$, and that the Brin-Katok Theorem 3.1 extends to local pressure:
$\int_{X}{\overline{P}}_{\mu}(f,\phi,x)\,d\mu=P_{\mu}(f,\phi):=h_{\mu}(f)+\int_{X}\phi\,d\mu.$
We have the following partial variational principle, extending Ma & Wen’s result Theorem 3.2.
@@
Given a continuous map $f:X\to X$ on a compact metric space $(X,d)$, a Borel subset $Z\subset X$ and a Borel probability measure $\mu$ on $X$, we have for any $s\geq 0$ and continuous $\phi:X\to{\mathbb{R}}$:
@@
$Z_{k}:=\left\{x\in Z:\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\leq s+\varepsilon\text{ for all }0<r<\frac{1}{3k}\right\}.$
@@
$\mu(B_{n_{i}(x_{i})}(x_{i},r))\geq\exp\left(-(s+\varepsilon)n_{i}(x_{i})+\sum_{m=0}^{{n_{i}(x_{i})}-1}\phi(f^{m}(x_{i}))\right).$
@@
$V_{k}:=\left\{x\in Z:\liminf_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\geq s-\varepsilon,\text{ for every }0<r<\frac{1}{k}\right\}.$
@@
$V_{k_{0},N}:=\left\{x\in Z:\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(x))-\log(\mu(B_{n}(x,r)))\right)\geq s-\varepsilon,\text{ for all }n\geq N,0<r<\frac{1}{k_{0}}\right\}.$
@@
Passing to the limit $N\to\infty$, we get $M_{Z}(s-\varepsilon,r,\phi)>0$. This means that $M_{Z}(r,\phi)>s-\varepsilon$. Finally, passing to the limit $r\to 0$ and taking into account that $\varepsilon$ is arbitrarily small, we obtain $P_{Z}(f,\phi)\geq s$. The proof is complete. $\Box$
@@ 3.4 The translocal version of measure-theoretic local pressure
In this section we introduce a translocal version of the upper measure-theoretic local pressure, which is a modification of ${\overline{P}}_{\mu}(f,\phi,z)$ from (9), in the following way.
$\overline{P}_{\mu,\omega}(f,\phi,z):=\limsup_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(z))-\log\mu(B(z,e^{-\omega n}))\right),$
@@
$\underline{P}_{\mu,\omega}(f,\phi,z):=\liminf_{n\to\infty}\frac{1}{n}\left(\sum_{m=0}^{n-1}\phi(f^{m}(z))-\log\mu(B(z,e^{-\omega n}))\right).$
@@
where the infimum is over all finite or countable covers of $Z$ by metric balls $\{B(x_{j},e^{-\omega n_{j}})\}_{j\in J}$ with $n_{j}\geq N$. Let
@@
The graph of the function $s\mapsto\overline{M}_{f,Z,\omega}(s,\phi)$ has a unique critical parameter $s_{0}$, where the function $s\mapsto\overline{M}_{f,Z,\omega}(s,r,\phi)$ drops from $\infty$ to 0. The graph of the function $s\mapsto\underline{M}_{f,Z,\omega}(s,r,\phi)$ has a unique critical parameter $s_{1}$, where the function $s\mapsto\underline{M}_{f,Z,\omega}(s,\phi)$ drops from $\infty$ to 0.
@@
A metric space $(X,d)$ is called boundedly compact if all bounded closed subsets of X are compact. In particular ${\mathbb{R}}^{n}$ and Riemannian manifolds (see [26, p.9]) are boundedly compact.
@@
(Vitali Covering Lemma, Theorem 2.1 in [33]). Let $(X,d)$ be a boundedly compact metric space and a family of closed balls ${\mathcal{B}}=\{\overline{B(x,r)}:x\in X,r>0\}$ in X such that
@@
$\bigcup_{B(x,r)\in{\mathcal{B}}}B(x,r)\subset\bigcup_{i\in I}\overline{B(x_{i},5\cdot r_{i})}.$
@@
Given a continuous map $f:X\to X$ on a compact metric space $(X,d)$, a Borel subset $Z\subset X$ and a Borel probability measure $\mu$ on $X$, we have for any $s\geq 0$ and continuous $\phi:X\to{\mathbb{R}}$:
@@
2) If $\mu(Z)>0$ and ${\underline{P}}_{\mu,\omega}(f,\phi,x)\geq s$ for all $x\in Z$, then $\underline{P}_{Z,\omega}(f,\phi)\geq s$.
@@
where the last inequality follows by the disjointness of metric balls in the family ${\mathcal{G}}_{N}$. Taking the $\limsup$ as $N\to\infty$, we get
@@
${\mathcal{C}}:=\left\{B(y_{j},e^{-\omega n_{j}}):y_{j}\in V_{N_{0}},~n_{j}\geq N_{0},~B(y_{j},e^{-\omega n_{j}})\cap V_{N_{0}}\neq\emptyset,j\in J\right\}.$
@@
Taking the $\liminf$ as $N\to\infty$, we get $\underline{M}_{Z,\omega}(s-\varepsilon,\phi)>0$. This means that $\underline{M}_{Z,\omega}(\phi)>s-\varepsilon$. Finally taking into account that $\varepsilon$ is arbitrarily small, we obtain $\underline{P}_{Z,\omega}(f,\phi)\geq s$. The proof is complete. $\Box$
@@ 4 Discussion on local entropy
In the literature on dynamical systems there are several attempts to define local entropy. Bowen [10] introduced the notion of $\varepsilon$-entropy to get an upper estimate of the difference between $h_{\mu}(f)$ and the entropy of a partition with diameter less than $\varepsilon$. Another notion of local entropy (called conditional topological entropy) was given by Misiurewicz [34], who proved that vanishing of his local entropy implies existence of maximal measure. Finding (uniqueness of) measures of maximal entropy was also the reason for Newhouse [35] and Buzzi & Ruette [15, 16] to define a version the local entropy $h_{\text{\tiny loc}}(f)$, see (6). Buzzi & Ruette [16] proved that for $C^{1}$ maps $f:[0,1]\to[0,1]$ with critical set $\operatorname{Crit}(f)=\{x\in[0,1]:f|_{U}\text{ is not monotone on each neighborhood }U\owns x\}$, if $h_{\text{\tiny top}}(f)>h_{\text{\tiny top}}(f,\operatorname{Crit}(f))+h_{\text{\tiny loc}}(f)$, then $f$ admits a measure of maximal entropy.
4.1 Orbit complexity
Brudno [13] introduced the notion of orbit complexity for a homeomorphism $f:X\to X$ of a compact space $X$ and related it to measure-theoretic entropy and topological entropy. In Brudno’s approach the orbit complexity is a measure of the amount of information that is necessary to describe the orbit. The complexity $K(x,f)$ of the orbit of $x\in X$ is defined in the spirit of asymptotically optimal functions in the sense of Kolmogorov; it indicates the limit proportion of $n$-prefixes of the symbolic itinerary of $x$ needed to reconstruct the whole $n$-prefix. In his Theorem 3.1, which slightly predates what we state as the Brin-Katok Theorem 3.1 in Section 3.1, Brudno proved the following result: If an $f$-invariant measure $\mu$ is ergodic, then [^2] Entropy is computed with respect to logarithmic base $b$ where $b$ is the cardinality of the alphabet used for the symbolic itineraries. $K(x,f)=h_{\mu}(f)$ for $\mu$-a.e. $x\in X$. Theorem 3.2 in [13] says that for any point $x\in X$ one has $K(x,f)\leq h_{f}(x)$, where $h_{f}(x):=\sup\{h_{\mu}(f):\mu\text{ is a weak accumulation point of }(\frac{1}{n}\sum_{k=0}^{n-1}\delta_{f^{k}(x)})_{n\geq 1}\}$ is the entropy of Kamae introduced in [29].
The results of Brudno were extended by White [47], who proved that for any ergodic $f$-invariant measure $\mu$, $\mu$-a.e. orbit has a limiting complexity equal to $h_{\text{\tiny top}}(f)$ Also, if $f:X\to X$ is minimal, then the set of points with orbit upper complexity equal to the topological entropy is a residual set. If $f:X\to X$ has a unique measure $\mu_{0}$ of maximal entropy, then any point with lower complexity equal to $h_{\text{\tiny top}}(f)$, is generic for the measure $\mu_{0}$. This is likely to hold also if each measure of maximal entropy is fully supported as in Corollary 2.1.
Another extension of results of Brudno [13] one can find in the paper by Galatolo [23], who considered a dynamical system given by a continuous map on a separable metric space, and introduced a notion of computable structure on a metric space with a new notion of orbit complexity with respect to this structure. For compact spaces, Galatolo’s definition coincides with the definition given by Brudno.
A different way of using single orbits to estimate entropy goes back to time series analysis, and is known as Takens’ estimator in applied dynamics and fractal dimension estimation, see e.g. [40, 43]. Viewing orbits $Y:=(f^{n}(x_{0}))_{n\geq 0}$ for $x_{0}\in X$ as time series, one can define $k\times\dim(X)$-dimensional reconstruction vectors $(x_{n},x_{n+1},\dots,x_{n+k-1})$ and estimate the probability ${\mathbb{P}}_{\varepsilon}^{k}(Y)$ of finding two such vectors at mutual distance $\leq\varepsilon$. The entropy is the logarithmic growth rate of this quantity as $\varepsilon\to 0$: $h(Y)=\lim_{\varepsilon\to 0}\lim_{k\to\infty}\frac{1}{k}\log{\mathbb{P}}^{k}_{\varepsilon}(Y)$. Takens’ estimator is a maximum likehood estimator of this quantity, which is also applicable to arbitrary time series, i.e., not just for orbits only.
4.2 Entropy pairs and tuples
In 1993, Blanchard [6] introduced the notion of entropy pairs, in order to localize “where” in the system entropy is generated. This was generalized to entropy $n$-tuples and entropy sets by Huang & Ye [28], Dou, Ye & Zhang [19] and Blanchard & Huang [7].